Starting a race , 68.2 kgsprinter runs the first 7.04 m in 1.60 sec, starting from the rest  and accelerating uniformly.
(A)what is the sprinter speed at the end of 1.60s?
(B)what is the sprinter k.E . ?
(C)what is the average power does the sprinter generate during the 1.60 s unterval?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
 
(A)
  s=0.5 a t^2
 7.04=0.5 a 1.6^2
 a=5.5
 
v=at
 
v=5.5*1.6
.
v=8.8 m/s
 
 
(B)
    K.E=0.5mv^2
   K.E=0.5*68.2*(5.5*t)^2
   K.E=11346.775 * t^2 J
          at any time t
 
(C)
  P=m*a*s/t
 P=68.2*5.5*7.04/1.6
P=1650.44 J/s
 
 
 
 
Regards
Arun (askIITians Forum expert)